find the surface area of the portion of the curve x^2+y^2=4 from x=0 to x=2 when it is revolved about the y-axis


1
Expert's answer
2022-05-30T23:07:53-0400
"S=2\pi\displaystyle\int_{a}^{b}r\cos (t)\sqrt{(r(t))^2+(r'(t))^2}dt"

"x^2+y^2=4"

"r=2, a=-\pi/2, b=\pi/2"

"S=2\pi\displaystyle\int_{-\pi/2}^{\pi/2}2\cos t\sqrt{(2)^2+(0)^2}dt"

"=8\pi[\sin t]\begin{matrix}
 \pi/2 \\
 -\pi/2
\end{matrix}=16\pi({units}^2)"